### 19 Jan 2024 v2

# Let’s do the time-warp-attend: Learning topological invariants of dynamical systems

Abstract
Dynamical systems across the sciences, from electrical circuits to ecological networks, undergo qualitative and often catastrophic changes in behavior, called bifurcations, when their underlying parameters cross a threshold. Existing methods predict oncoming catastrophes in individual systems but are primarily time-series-based and struggle both to categorize qualitative dynamical regimes across diverse systems and to generalize to real data. To address this challenge, we propose a data-driven, physically-informed deep-learning framework for classifying dynamical regimes and characterizing bifurcation boundaries based on the extraction of topologically invariant features. We focus on the paradigmatic case of the supercritical Hopf bifurcation, which is used to model periodic dynamics across a wide range of applications. Our convolutional attention method is trained with data augmentations that encourage the learning of topological invariants which can be used to detect bifurcation boundaries in unseen systems and to design models of biological systems like oscillatory gene regulatory networks. We further demonstrate our method’s use in analyzing real data by recovering distinct proliferation and differentiation dynamics along pancreatic endocrinogenesis trajectory in gene expression space based on single-cell data. Our method provides valuable insights into the qualitative, long-term behavior of a wide range of dynamical systems, and can detect bifurcations or catastrophic transitions in large-scale physical and biological systems.
Keywords: dynamical systems, bifurcations, topological invariance, Hopf bifurcation, physics-informed machine learning, augmentation, single-cell RNA-sequencing.
1 Introduction
Gradually changing the parameters of a dynamical system will typically result in innocuous, quantitative alterations to the system’s behavior. However, in exceptional cases, parameter perturbations will result in a qualitative, potentially catastrophic change in the system’s dynamical behavior called a bifurcation. Bifurcations may lead to extreme consequences, as when a beam buckles and breaks under increasing weight (Strogatz, 1994), when disastrous oscillations presage the collapse of poorly designed bridges, like the infamous Tacoma Narrows Bridge (Billah & Scanlan, 1991; Deng et al., 2016), or when an airplane exceeds a speed threshold, leading to dangerous self-excited oscillations in its tail (Weisshaar, 2012; Dimitriadis, 2017). In biological systems, bifurcations marking the transition away from oscillatory dynamics on cellular or organ-level scale, such as those regulating blood pumping and circadian rhythms, can have dire health consequences (Xiong & Garfinkel, 2023; Qu et al., 2014).
Understanding bifurcations, be they in physics, engineering, or biology, is a subject of intense research in the study of nonlinear systems (Kuznetsov, 1998; Karniadakis et al., 2021; Ghadami & Epureanu, 2022). The central challenge across all applications is that bifurcations can be qualitatively identical (e.g. the emergence of oscillations) while being quantitatively very different (e.g. the amplitude or frequency of those oscillations); smoothly warping the state space has no qualitative effect on the dynamics. Put another way, a bifurcation detector must be selective for topological structure of underlying dynamics but invariant to the geometric nuisances of those dynamics. Predicting the onset of these bifurcations often requires the explicit use of governing equations, which must be hardwired or inferred. Hence, the analysis of real data, where such equations are often unknown or difficult to define, is particularly challenging (Kuznetsov, 1998). Applications to real data either use hardwired (Wissel, 1984; Wiesenfeld, 1985; Scheffer et al., 2009) or learned (Bury et al., 2023) time series features, but no approach, to our knowledge, explicitly addresses the selectivity-invariance trade-off inherent to the topological nature of the problem. Additionally, where vector field data is readily available and time-series tracing is not, e.g., in capturing gene expression of cells in high-throughput single-cell experiments, it remains unclear how to extract topological information either from such noisy and sparse vector fields directly or from their integration.
We address this challenge by learning a universal dictionary of dynamical features which is explicitly encouraged to be invariant to geometric nuisances but selective for the topological features which define a particular type of bifurcation (Fig. 1). This is achieved primarily by the use of topological data augmentation: each sample of our training data is a velocity field which comes bundled with numerous deformed but identically labeled versions. The features extracted from these augmented samples are then passed to a supervised classifier which must contend with the substantial geometric nuisances induced by the augmentation scheme. We demonstrate our method for the ubiquitous supercritical Hopf bifurcation whereby a stable fixed point (or “equilibrium”) loses stability and transitions into a regime exhibiting a stable limit cycle. We train on synthetic, augmented data generated from a simple oscillator prototype. Despite being only trained on data from one simple equation, our architecture can generalize to radically different systems, such as the nonlinear Selḱov glycolysis oscillator and the Belousov-Zhabotinsky model of chemical oscillations, precisely because of our use of topological augmentation.
Our approach is geared towards realistic, high-dimensional, noisy, and complex data. For instance, we apply our technique to the repressilator, a gene regulatory network which can generate oscillatory gene expression, and infer the probability for oscillatory behavior under conditions of varying gene transcription and degradation rates. Finally, we apply our method to the analysis of complex, large-scale biological data. By extracting features from single-cell gene expression velocity data during the differentiation of pancreatic cells, our method can distinguish between proliferation and differentiation dynamics with high fidelity. Our results indicate that our approach can be applied to a broad range of dynamical systems, providing valuable insights into their dynamic characteristics and bifurcation behavior  [^1] Code available at: https://github.com/nitzanlab/time-warp-attend.
Figure 1: Overview. Our framework classifies long-term dynamical behavior of point vs periodic attractor across systems and regimes by leveraging topologically-invariant augmentations of a single prototypical system. To encourage generalization, we (1) abstract system-specific vector magnitudes and use angular representations instead, and (2) deploy attention layers which focus on learning essential dynamical cues. Using this framework, we are able to classify and recover the bifurcation boundaries of complex, unseen systems. We apply this method to the problem of recovering cell-cycle scores which distinguish between proliferating and differentiating cell populations. All vector fields are plotted as streamlines.
The main contributions of this paper are:
  1. 1.
    A novel approach to topologically invariant feature learning in dynamical systems. The approach uses warped vector field data as augmentations to generalize across systems exhibiting similar dynamics in terms of their topological representation in phase space, allowing for a more selective and invariant understanding of bifurcations in diverse real-world systems.
  2. 2.
    Invariant classification of complex and nonlinear synthetic data, including models of diverse chemical and biological systems, achieved by leveraging knowledge of a single prototypical system.
  3. 3.
    The identification of bifurcation boundaries using classifier robustness.
  4. 4.
    The recovery of distinct proliferation and differentiation dynamics along the pancreatic differentiation trajectory from single-cell gene expression data, providing a practical application of our method to the analysis of biological systems.
2 Related Work
Classical methods for detecting limit cycles Solving non-linear differential equations is in general difficult, and it has been recognized since the time of Poincaré that reasoning about dynamical systems’ characteristics can be made easier by qualitatively delineating solutions according to their long-term behaviors. For autonomous (i.e., time-invariant) systems in two dimensions, if the dynamics do not diverge, these behaviors are either fixed points or limit cycles (see Methods), and transitions between these dynamical regimes as a result of changing parameters are called bifurcations. More explicitly, a bifurcation corresponds to a change in the topological structure of a phase portrait as a parameter of the dynamical system is varied. Full characterization of the location and number of limit cycles in two-dimensional, polynomial systems is a challenging open problem in dynamical systems theory (Hilbert, 2019; Smale, 2000).
Early theoretical results guaranteeing the existence or absence of a limit cycle in two-dimensional autonomous systems are the theorems of Poincaré-Bendixson and Bendixson-Dulac, respectively. Further guarantees are possible when the system can be reduced to the form of a so-called Liénard equation. None of these results, however, is constructive or easy to use in practice. Center manifold theory (Carr, 2012), for its part, analyzes dynamics on representative subspaces where systems reduce to normal forms having known bifurcation behaviors, though this method requires knowledge of the underlying governing equation. The same limitation applies to Conley-Morse theory, which represents dynamics as a graph whose nodes correspond to long-term behaviors (Conley, 1978). Importantly, none of these approaches is adapted to real data, where equations are unknown and system measurements can be sparse, incomplete and noisy.
Data-driven methods for detecting limit cycles Data-driven methods fall into non-topological and topological camps. In the former camp, several approaches use compressed sensing of trajectory data to reduce noise, followed by identification of system parameters and associated bifurcation diagrams (Wang et al., 2011; 2016), though this method requires inferring governing equations for each new sample. Other work sought to reduce vector field measurements to a Conley-Morse graph (Chen et al., 2007b; a) for which a database over parameter regimes can be constructed via combinatorial dynamics exploration and graph equivalence comparisons (Arai et al., 2009). Yet other approaches used estimates of Lyapunov exponents to delineate between periodic and stationary systems, although, as we show below, this approach can be sensitive to noise (Stoker, 1950; Zhang et al., 2009).
On the topological front, there is a growing body of work using persistent homology to detect periodicity in time series data (Ravishanker & Chen, 2021). For example, Perea & Harer (2015) extracted persistence barcodes from delay embeddings of time series to measure periodicity, although this process is very computationally expensive and there is no standard approach to classifying barcodes. Persistent homology approaches often assume that the underlying bifurcation parameter is known (Tymochko et al., 2020), which greatly simplifies the problem at the cost of realism. To our knowledge, there is no data-driven method which is efficient (not requiring expensive barcode calculations for each sample), general (applicable to novel systems without retraining) and robust (not sensitive to noise).
Generalization across dynamical systems. Comparatively few approaches exist for the simultaneous analysis of diverse dynamical systems and their bifurcation behavior. Work by Brown & Sethna (2003); Transtrum & Qiu (2014) and Quinn et al. (2021) mapped families of dynamical systems to a manifold whose geometry encodes different qualitative behaviors, but these methods have no straightforward way of identifying bifurcations between these behaviors. More recent work from Ricci et al. (2022) uses a data-driven approach called phase2vec to learn manifolds of dynamical systems, leading to dynamical embeddings which encode qualitative properties of the underlying systems, though it is not clear if this method can detect or distinguish dynamical regimes and their associated bifurcations per se.
Diffeomorphic augmentations for generalization. Diffeomorphic transformations which smoothly preserve topology in the context of dynamical systems have been used principally in computer vision. In order to train a classifier of a limited image set of biological entities such as diatoms, Vallez et al. (2022) used diffeomorphic augmentations to expand its train data, however, since the labels of these data were not necessarily preserved by these transformations, a manual post-filtering was required to maintain realism. In a second set of works concerning dynamical systems specifically, Khadivar et al. (2020) and Bevanda et al. (2022) fit a diffeomorphic transformation into closed or linearized forms, and work by Bramburger et al. (2021) used an autoencoder approach to reduce a complex dynamical system to a topologically conjugate (i.e., dynamically equivalent) form of discrete-time mapping. Here, we use diffeomorphisms to curate a set of consistently labeled dynamical augmentations, see Appendix 7.1.1.
3 Methodology
Topological equivalence in dynamical systems Let $\dot{x}=f(x;\theta)$ for $x\in\mathbb{R}^{2}$ be a parameterized dynamical system with parameter vector $\theta\in\Theta$. A solution to this equation, $F(x_{0},t)$, with an initial condition, $x(t=0)=x_{0}$, is called a trajectory through $x_{0}$. A standard approach to understanding dynamical systems (Kuznetsov, 1998) is to characterize them by the topology of all of their trajectories, construed as functions of $x$ and $t$ or just $x$. For example, do trajectories extend indefinitely, stop at a point (a fixed point) or loop back on themselves (a limit cycle)? More precisely, we say that $f$ has a fixed point at $x^{*}$ and $\theta=\theta^{*}$ if $f(x^{*};\theta^{*})=0$. We say that $f$ has a limit cycle at $\theta^{*}$ if there exists a trajectory of $f$ which is closed and isolated (neighboring trajectories are not closed). Importantly, these properties are preserved when the state space is subjected to a topology-preserving map, making them topological invariants of the dynamical systems. However, these properties may change as $\theta$ changes; i.e. the system undergoes a bifurcation. For example, in a system that undergoes a supercritical Hopf bifurcation, there exists a parameter $\theta$ such that below a certain threshold of $\theta$ the system exhibits a stable fixed point, and above it, the fixed point loses its stability and a stable limit cycle emerges.
Dynamical equivalence and topological augmentation
In what follows, we will consider autonomous systems with either a single stable fixed point (i.e., equilibrium) or a stable limit cycle with an unstable fixed point in its interior. We refer to systems, $f$, in one of these two categories as belonging to the same dynamical class, denoted $y(f)$, and we denote the fixed point and limit cycle classes considered here the “pre-Hopf” and “post-Hopf” classes, respectively. As discussed above, these classes are invariant to topology-preserving deformations of the state space. Here, we will use diffeomorphisms, $h\in H$ which induce an equivalence relation on dynamical systems in a standard and well-studied way: we have $y(f)=y(f\circ h)$ for all $h\in H$ so that a given $H$ defines an equivalence relation which partitions the space of dynamical systems into its dynamical classes. We refer to two systems, $f,g$ such that $g=f\circ h$ as being dynamically equivalent  [^2] We choose the neutral term “dynamical equivalence” to distinguish our notion from the similar notions of topological conjugacy and topological equivalence (Kuznetsov, 1998). See 7.1.1 for details..
In this light, classification becomes a problem of learning representations of dynamical systems which are invariant to all $h\in H$ but selective for the underlying dynamical class. To that end, we learn features on numerous, transformed versions of a prototypical system, a simple oscillator $f$ defined for a point at radius $r$ and angle $\phi$ with parameters $\theta=(a,\omega)$:
$\dot{r}$ $=r(a-r^{2})$
$\dot{\phi}$ $=\omega,$
which has a sparse functional form and undergoes a supercritical Hopf bifurcation for any $\omega\neq 0$ at $\theta^{*}=(0,\omega)$. We validate these learned features by measuring how well they encode the dynamical class and bifurcation boundary of testing data, which, though superficially very different from the training data, having different functional forms and parameter regimes, are nevertheless in the same equivalence classes induced by $H$.
Our training data is constructed by first acquiring planar vector fields, $\{f(x,\theta)\in\mathbb{R}^{2}\}$ evaluated at all points, $x$, on a fixed, $64\times 64$ lattice. Parameters, $\theta$, are sampled uniformly on $a\in[-0.5,0.5],\omega\in[-1,1]$, and each system is assigned a label according to whether it is in the pre- or post-Hopf dynamical class. Then, the dataset is augmented with expressive diffeomorphisms generated using monotonic rational-quadratic splines. In particular, we use the spline construction from Durkan et al. (2019), where instead of training the neural spline flows proposed in that work, we use its random initialization which generates a diverse set of diffeomorphic augmentations. For details on data augmentation, see Appendix 7.1.2.
Convolutional attention training. We construe our full, augmented data set as a collection of labeled arrays in $\mathbb{R}^{2\times 64\times 64}$ which we use to train a convolutional network. Our architecture is adapted from the discriminator of self-attention generative adversarial network (SAGAN) of Zhang et al. (2019), coupled with a multi-layered perception (MLP). For standardization of the data, we reduce each vector field to only its angular components, ignoring magnitudes which carry little information about cyclic behavior. The feature extraction module consists of four convolutional blocks with self-attention mechanism into the last two convolutional layers. Our reasoning for incorporating self-attention is that only a few dynamical keypoints in the vector field (i.e., the fixed point or the limit cycle) carry information about class label, and the rest of the array can be (nearly) ignored. Together, the data warping and attentional components of our pipeline comprise our eponymous method. Features from this network are passed to a multilayer perceptron classifier which is trained with a supervised, cross-entropy loss. Because of our augmentation scheme, this classifier must contend with data which, though radically different in appearance, belong to the same dynamical class, encouraging the classifier to learn the true dynamical invariances underlying the Hopf bifurcation. Training and architecture details are described in the Appendix.
4 Results
4.1 Transferring knowledge of dynamical classes from simple prototypes to complex test data
We evaluate our model and its baseline competitors on a diverse, synthetic dataset of noisy (zero-mean gaussian), real-world systems drawn from across the sciences. Test datasets include the simple harmonic oscillator (“SO”), its warped variant (“Augmented SO”), a system undergoing a supercritical Hopf bifurcation as described in (“Supercritical Hopf”), examples based on Liénard equations (“Liénard Polynomial” and “Liénard Sigmoid”), and systems used to model oscillatory phenomena in nature and engineering applications, namely the van der Pol system, used for modeling electrical circuits in early radios (“Van der Pol”), the Belousov-Zhabotinsky model, a model of chemical oscillations (“BZ reaction”), and the Selḱov oscillator, used in modeling glycolitic cycles (“Selḱov”), see Appendix 7.2 and Strogatz (1994).
Our baselines comprise several competing feature types: (1) “Critical Points”, a heuristic based on critical points identification as derived from Helman & Hesselink (1989; 1991) algorithm, (2) “Lyapunov” which distinguishes pre- and post-Hopf states according to the value of the Lyapunov exponent of a time series integrated from the vector field, (3) “Phase2vec”, vector field embeddings from Ricci et al. (2022), a state-of-the-art deep-learning method for dynamical systems embedding, and (4) “Autoencoder”, latent features of of a convolutional vector field autoencoder (the same as that used in phase2vec training), see Appendix 7.3. As with our method, for all baselines except Critical Points, classifications are produced by a linear classifier trained on features from augmented simple harmonic oscillator data (Augmented SO) in order to discriminate between point and cyclic attractors. The accuracy average and standard deviation for our model and the trained baselines are computed over 50 training runs.
We find that our framework generalizes across a wide range of dynamical systems and classifies distinct dynamical regimes in noisy (zero-mean gaussian) conditions better on average than baseline methods (see Table A4, Table 1 and examples at Fig. A8). Critical Points achieves nearly perfect classification on noiseless data but is easily disrupted by noise (e.g., the perfect classification of Supercritical Hopf drops to 56% with Gaussian noise $\sigma=0.1$). Likewise, Lyapunov suffers a substantial decrease in accuracy for classical oscillators (e.g., 89% accuracy of Supercritical Hopf reduces to 70% accuracy with Gaussian noise $\sigma=0.1$). Both with and without noise, Phase2vec extends mainly to SO and Supercritical Hopf. Autoencoder obtains competitive results, however, it exhibits little generalization to Liénard Polynomial system, and essentially no generalization to Selḱov system. We also test resilience to noise in the Augmented SO data (see Table 2). Despite the substantial distortion of the vector fields caused by noise, the accuracies of our model and that of autoencoder representation only gradually decrease, unlike the sharp decrease in accuracy for the other baselines.
Table 1: Test accuracy with added Gaussian noise of $\sigma=0.1$.
SO Aug. SO Supercritical Hopf Liénard Poly Liénard Sigmoid Van der Pol BZ Reaction Selḱov Average
Our Model 0.98$\pm$0.01 0.93$\pm$0.01 0.99$\pm$0.01 0.86$\pm$0.13 0.92$\pm$0.07 0.83$\pm$0.15 0.82$\pm$0.11 0.65$\pm$0.03 0.87
Critical Points 0.54 0.56 0.56 0.70 0.49 0.56 0.84 0.51 0.60
Lyapunov 0.64 0.60 0.71 0.85 0.67 0.87 0.83 0.57 0.72
Phase2vec 0.73$\pm$0.08 0.71$\pm$0.04 0.88$\pm$0.03 0.49$\pm$0.06 0.48$\pm$0.0 0.49$\pm$0.04 0.5$\pm$0.0 0.49$\pm$0.0 0.6
Autoencoder 0.95$\pm$0.03 0.87$\pm$0.01 0.97$\pm$0.01 0.63$\pm$0.16 0.88$\pm$0.09 0.81$\pm$0.16 0.82$\pm$0.13 0.49$\pm$0.02 0.8
Table 2: Evaluation of Augmented SO with added Gaussian noise of scale $\sigma$.
$\sigma$ 0 0.10 0.20 0.30 0.40 0.50 0.60 0.70 0.80 0.90 1.00
Our Model 0.93$\pm$0.01 0.93$\pm$0.01 0.93$\pm$0.01 0.91$\pm$0.02 0.86$\pm$0.04 0.78$\pm$0.06 0.7$\pm$0.06 0.64$\pm$0.04 0.6$\pm$0.04 0.57$\pm$0.03 0.54$\pm$0.02
Critical Points 0.93 0.59 0.53 0.507 0.51 0.51 0.51 0.51 0.51 0.50 0.50
Lyapunov 0.71 0.60 0.52 0.50 0.50 0.50 0.50 0.50 0.50 0.50 0.50
Phase2vec 0.87$\pm$0.03 0.7$\pm$0.05 0.53$\pm$0.01 0.51$\pm$0.01 0.5$\pm$0.0 0.5$\pm$0.0 0.5$\pm$0.0 0.5$\pm$0.0 0.5$\pm$0.0 0.5$\pm$0.0 0.5$\pm$0.0
Autoencoder 0.86$\pm$0.01 0.87$\pm$0.01 0.84$\pm$0.02 0.8$\pm$0.04 0.76$\pm$0.05 0.72$\pm$0.05 0.69$\pm$0.06 0.68$\pm$0.06 0.65$\pm$0.06 0.64$\pm$0.05 0.62$\pm$0.05
In the Appendix 7.3, we report the results of another baseline, “Parameters” which trains a linear classifier based on a flattened vector of coefficients of a degree three polynomial fit to each vector field with least squares, a common method for bifurcation analysis, although this resulted in a poor average accuracy of 53%.
Furthermore, we find that all modules in our pipeline (topological augmentation, angular representation of data, self-attention layers) contribute, on average, to accuracy in complex systems, see Appendix 7.4. Concretely, the average classification accuracy of the framework, trained on data without augmentations, using vector representation, lacking attention layers, or without any of these, is 78%, 82%, 77%, 68%, respectively, and notably, except for the ablation of attention layers, other ablated models do not generalize to the complex BZ reaction and Selḱov systems. We also found that topological augmentations are particularly necessary when generalizing to systems with a fixed point far from the center of the grid. For example, “centered” systems such as Liénard Sigmoid and Liénard Polynomial, are well-characterized without augmentations, reaching 0.96 and 0.94 accuracy respectively, whereas systems that are not centered, e.g. BZ reaction whose fixed point is close to the edge of the grid, require augmentations for generalization.
We also evaluate the robustness of our method as a function of (1) the training dataset size, where we find good generalization even on relatively small data sets (0.84 accuracy after training on $\sim 2k$ vector fields), (2) vector field resolution, where performance is stable when doubling the vector resolution (0.86 accuracy), and (3) architectural and optimization hyperparameters (learning rate, training epochs, etc.), obtaining almost consistently an overall accuracy of between 0.85 and 0.88. See Appendix 7.4 and Table A4 for details.
4.2 Recovering bifurcation boundaries
Classifier confidence is expected to decrease for systems near the bifurcation boundary since limit cycles in this regime shrink below the vector field resolution, making them indistinguishable from fixed points. This effect can be used to pinpoint the bifurcation boundary in parameter space. For example, when we plot each system’s average prediction as a function of its bifurcation parameter (for Augmented SO, the parameter values are chosen as the original $a,\omega$), we find an inverse relation between the parametric distance from the bifurcation boundary and classifier confidence (see Fig. 2, Fig. A11 and Fig. A12 for bifurcation diagrams of alternative baselines). An exception is the upper boundary of the Selḱov system where instances of high $b$ value are misclassified as exhibiting cyclic dynamics, and hence, the upper bifurcation boundary in the $a-b$ parameter plane is undetected (Fig. 2, right plot). We explain the systematic misclassification as a result of damping which makes trajectories move in slow, tight spirals towards the fixed point attractor in a manner numerically indistinguishable from limit cycles (Fig. A7). These overall suggest that our method can be geared towards recovering the bifurcation boundaries of unseen systems.
Figure 2: Inference of bifurcation boundary. Each system is plotted at coordinates of its simulated parameter values and colored according to its average cycle prediction across 50 training re-runs. The true parametric bifurcation boundary is plotted in black. Colors are scaled between a perfect cycle classification (red) and perfect point classification (blue) with white at evenly split classification.
4.3 Predicting oscillatory regulation of the repressilator system
In order to demonstrate the viability of our method for higher-dimensional systems under realistic noise conditions, we apply our topological approach to a computational model in a well-studied biological setting, the so-called repressilator model (Elowitz & Leibler, 2000). The repressilator system is a synthetic, genetic regulatory network that was introduced by Elowitz and Liebler into bacteria to induce oscillatory regulation of gene expression Fig. 3. The system is modeled by the mRNA and protein counts of three genes (TetR, LacI and $\lambda$ phage cI) that inhibit each other in a cyclic manner (see Appendix 7.7 for governing equations). Theoretically and experimentally, the existence and characteristics of oscillations have been shown to depend on numerous factors like the rates of transcription, translation, and degradation, as well as intrinsic and extrinsic noise (Verdugo, 2018; Potvin-Trottier et al., 2016). Here we test our method in classifying a regime of the repressilator which has been shown to be undergoing a Hopf bifurcation, though, unlike the previous complex simulated systems, it spans six dimensions and is constructed from noisy cellular time-series.
To generate 2D vector field data, we consider trajectories of cellular gene expression measured by protein levels of TetR and LacI (as in Fig. 3, middle). Different instances of simulations share the initial condition but vary by the rate of transcription, $\alpha$, and by the ratio of protein and mRNA degradation rates, $\beta$. We sample 100 cell states along the temporal trajectory and add Gaussian noise of $\sigma=0.5$ to the gene expression states (to protein and mRNA levels) to model intrinsic noise. For each set of measurements, we interpolate an input 64x64 vector field, based on their velocities across TetR-LacI protein phase space. For further details of the data generation process see Appendix 7.7. We predict for each vector field its dynamical class, achieving 87% classification accuracy and proximate bifurcation boundary to the analytically-derived boundary (Fig. 3, right (Verdugo, 2018)). In Fig. A13 we show prototypes of trajectories across the $\alpha-\beta$ parameter space, where we see that where we fail to classify a point attractor, trajectories of cells loop closely in the TetR-LacI protein plane and generate data samples that highly resemble trajectories with a periodic attractor rather than with a point attractor.
Figure 3: Oscillations in repressilator system. Example of oscillations in the repressilator system viewed as a time-series (left) or projected onto TetR-LacI protein plain (middle). (right) Predicted oscillatory behavior by our model output upon varying transcription rate without repression ($\alpha$) and the ratio of protein and mRNA degradation rates ($\beta$). The theoretical boundary is marked in black. See details in Appendix 7.7.
4.4 Distinguishing dynamical regimes in single-cell data of the pancreas
Last, we leverage our method to characterize distinct dynamical regimes corresponding to different biological processes in real data. Specifically, we focus on distinguishing proliferation vs. differentiation dynamics in single-cell RNA-sequencing data. A population of proliferating, precursor cells expresses genes associated with DNA duplication in a cyclic fashion (Tirosh et al., 2016; Schwabe et al., 2020; Karin et al., 2023). Cell differentiation, on the other hand, and similarly to the Waddington landscape abstraction (Waddington, 2014), is often characterized by cellular flow in gene expression space towards a localized, differentiated state (Sáez et al., 2022). We show that our model can detect the qualitative change in the flow of cells in gene expression space as they mature and differentiate out of their precursor cell state, moving from a dynamical regime characterized by a periodic behavior to that characterized by fixed points.
We focus on the process of pancreas endocrinogenesis where Ductal and Ngn3+ precursor cells rapidly proliferate to give rise to functional cell types Alpha, Beta, Delta and Epsilon (Bastidas-Ponce et al., 2019). We examine a single-cell RNA-sequencing dataset where the expression of over 2,000 genes are measured per cell, for a total of 3,600 cells. While the RNA expression profile, providing a gene expression vector per cell, provides for each cell its location within gene expression space, RNA velocity (Bergen et al., 2021; Gorin et al., 2022) can be used to compute the effective velocity of each cell (based on differences in spliced and unspliced RNA counts). We use scVelo (Bergen et al., 2020) to infer cellular velocities in the high-dimensional gene expression space and to project them to a 2D UMAP representation (Fig. 4, left). Simultaneously, we also compute for each cell a ‘cell-cycle score’, which captures the extent of expression of cell-cycle genes (in phase S and in grouped G2 and M phases, Fig. 4, middle; see Appendix 7.8, Bergen et al. (2020); Tirosh et al. (2016)) . For each set of cells, we interpolate an input 64x64 vector field, based on their RNA velocities, over the 2D UMAP representation.
To generate multiple frames from the trajectory, we randomly split cells of each cell type into groups of at least 50 cells resulting in 69 (sparse, noisy) vector fields with corresponding cell-cycle score (Fig. 4, middle, see details in Appendix 7.8). We then used our model (still trained solely on simulated augmented simple oscillator data) to characterize the dynamic regimes of the cellular behavior. Our model can successfully identify proliferative, cyclic behavior within the single-cell data, as reflected by 94% accuracy in predicting a cyclic behavior matching the high cell cycle score (Fig. 4, right), exceeding the prediction accuracy reached by alternative baselines (Critical Points: 54%, Lyapunov: 49%, Phase2vec: 30%, Autoencoder: 30%). As opposed to the cell-cycle score, our model does not use any prior knowledge regarding cell-cycle related genes, and is only based on the structural features of the cellular vector fields, and therefore can also be employed when such prior knowledge is not available.
Figure 4: Proliferation-to-differentiation transition in pancreas development. (left) Gene expression velocity of cells undergoing pancreas endocrinogenesis in UMAP space. Colors distinguish between different cell types. (middle) Cell-cycle score based on the fraction of cells with high S-phase gene expression (with S-score $>$ 0.4; see Appendix 7.8). (right) Predicted cell-cycle score computed as the average classification prediction across 50 training re-runs.
5 Limitations
Our framework learns to be invariant to geometric (as opposed to topological) properties of systems through topological data augmentation and convolutional attention, achieving competitive results in bifurcation detection across real-world systems. Nevertheless, we note here several important limitations of our method in its current form. The method is currently confined to the important but circumscribed setting of two-dimensional dynamics. Scaling up to much higher dimensions in these contexts presents technical challenges like the exponential growth of input size with dimension, as well as theoretical challenges such as the lack of structural stability in higher dimensions and the introduction of chaotic regimes. Since inherently low-dimensional dynamics often underlie high-dimensional measurements, especially in biology (Xiong & Garfinkel, 2023; Khona & Fiete, 2022), high dimensions can be mitigated by hardwired, or optimized dimensionality reduction methods as we deploy for the six-dimensional repressilator and for analyzing over 2,000 gene expression dynamics of the pancreas single-cell dataset. Additionally, we used a supervised, multi-class setup so that the classifier can predict that both a limit cycle and a fixed point co-exist. Empirically, we find that our method learns that the point and periodic attractors are exclusive, in fact, the learned logits consistently negate each other. For emphasis, we tested our trained model on samples representing a subcritical Hopf bifurcation having a regime with point and periodic attractors whose logits were of intermediate value but still negations, see Appendix 7.6. In the future, we aim to extend to combinatorial annotations of invariant sets, as elaborated in the discussion.
6 Discussion
The problem we consider is the data-driven classification of a diverse set of both real and synthetic dynamical systems into topological equivalence classes. This broad scope fundamentally distinguishes our work from most existing methods in bifurcation analysis which deal exclusively with the prediction of oncoming catastrophes from time series data of single, isolated systems. The fact that our method can be simultaneously applied to numerous systems of diverse functional and parametric forms demonstrates that our model has learned an abstract notion of these equivalence classes: it can be applied out-of-the-box to novel scenarios without the need for retraining. The integration of physics-informed data augmentation within the pipeline is crucial for attaining this high-level performance for complex systems that are substantially different from the prototype system underlying the training data. Our model captures visual uncertainty in systems near bifurcation boundaries, allowing for the extraction of an equation-free representation of these transitions. Furthermore, our model demonstrates robustness across samples disrupted by noise, withstands data produced from complex, high-dimensional time-series that is prominent for real-world applications such as the design of synthetic gene regulation, and successfully detects in real single-cell data the qualitative shift in cellular flow topology, corresponding to shifts of biological behaviors, from precursor to differentiated cell states, without relying on gene semantics.
Our work suggests numerous next steps including further translating our prior knowledge of topological properties, like the importance of critical points, into either explicit features or trainable, physics-informed network components (Karniadakis et al., 2021). Another goal is to broaden the scope to encompass a wider range of invariance types, combinatorial comparisons of invariant sets (e.g., via approximation of Conley-Morse graphs), and extensions to stronger notions of equivalence, such as topological conjugacy. To accommodate more versatile invariances and bifurcations, future work could extend our framework to the unsupervised case. For example, given that our data augmentations produce samples which are in the same dynamical equivalence class, a contrastive loss could be used to shape the feature space so that bifurcation classes are represented implicitly during training. Or, while challenging, following Skufca & Bollt (2008), a purely unsupervised approach could be used to explicitly construct warpings between systems in order to characterize equivalence classes.
Our work demonstrates the potential for scaling up to applications such as identifying dynamical regimes and underlying driving factors across both biological and physical applications.
Acknowledgments
We thank Yedid Hoshen for early discussions. This work was supported by the Israeli Council for Higher Education Ph.D. fellowship (N.M.), the Center for Interdisciplinary Data Science Research at the Hebrew University of Jerusalem (N.M.), the Zuckerman Postdoctoral Program (M.R.), Azrieli Foundation Early Career Faculty Fellowship, Alon Fellowship, and the European Union (ERC, DecodeSC, 101040660) (M.N.). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### 21 Mar 2024 v3

@@
 Our method provides valuable insights into the qualitative, long-term behavior of a wide range of dynamical systems, and can detect bifurcations or catastrophic transitions in large-scale physical and biological systems.
-Keywords: dynamical systems, bifurcations, topological invariance, Hopf bifurcation, physics-informed machine learning, augmentation, single-cell RNA-sequencing.
+Keywords: dynamical systems, bifurcations, topological invariance, Hopf bifurcation,
+physics-informed machine learning, augmentation, single-cell RNA-sequencing.
 1 Introduction
@@
 In biological systems, bifurcations marking the transition away from oscillatory dynamics on cellular or organ-level scale, such as those regulating blood pumping and circadian rhythms, can have dire health consequences (Xiong & Garfinkel, 2023; Qu et al., 2014).
-Understanding bifurcations, be they in physics, engineering, or biology, is a subject of intense research in the study of nonlinear systems (Kuznetsov, 1998; Karniadakis et al., 2021; Ghadami & Epureanu, 2022).
-The central challenge across all applications is that bifurcations can be qualitatively identical (e.g.
+Understanding bifurcations which transition between dynamical regimes, be they in physics, engineering, or biology, is a subject of intense research in the study of nonlinear systems (Kuznetsov, 1998; Karniadakis et al., 2021; Ghadami & Epureanu, 2022).
+One of the central challenges across all applications is that distinct dynamical regimes can be qualitatively identical (e.g.
 the emergence of oscillations) while being quantitatively very different (e.g.
 the amplitude or frequency of those oscillations); smoothly warping the state space has no qualitative effect on the dynamics.
-Put another way, a bifurcation detector must be selective for topological structure of underlying dynamics but invariant to the geometric nuisances of those dynamics.
-Predicting the onset of these bifurcations often requires the explicit use of governing equations, which must be hardwired or inferred.
+Put another way, a detector for different dynamical regimes and the transitions between them must be selective for topological structure of underlying dynamics but invariant to the geometric nuisances of those dynamics.
+Predicting the onset of such bifurcations often requires the explicit use of governing equations, which must be hardwired or inferred.
 Hence, the analysis of real data, where such equations are often unknown or difficult to define, is particularly challenging (Kuznetsov, 1998).
@@
 We train on synthetic, augmented data generated from a simple oscillator prototype.
-Despite being only trained on data from one simple equation, our architecture can generalize to radically different systems, such as the nonlinear Selḱov glycolysis oscillator and the Belousov-Zhabotinsky model of chemical oscillations, precisely because of our use of topological augmentation.
+Despite being only trained on data from one simple equation, our architecture can generalize to radically different systems, such as the nonlinear Selḱov glycolysis oscillator and the Belousov-Zhabotinsky model of chemical oscillations.
 Our approach is geared towards realistic, high-dimensional, noisy, and complex data.
@@
 2 Related Work
-Classical methods for detecting limit cycles Solving non-linear differential equations is in general difficult, and it has been recognized since the time of Poincaré that reasoning about dynamical systems’ characteristics can be made easier by qualitatively delineating solutions according to their long-term behaviors.
-For autonomous (i.e., time-invariant) systems in two dimensions, if the dynamics do not diverge, these behaviors are either fixed points or limit cycles (see Methods), and transitions between these dynamical regimes as a result of changing parameters are called bifurcations.
-More explicitly, a bifurcation corresponds to a change in the topological structure of a phase portrait as a parameter of the dynamical system is varied.
+Classical methods for detecting limit cycles.
+Solving non-linear differential equations is in general difficult, and it has been recognized since the time of Poincaré that reasoning about dynamical systems’ characteristics can be made easier by qualitatively delineating solutions according to their long-term behaviors.
+For autonomous (i.e., time-invariant) systems in two dimensions, if the dynamics do not diverge, these behaviors are either fixed points or limit cycles, and transitions between these dynamical regimes as a result of changing parameters are called bifurcations.
+More explicitly, a bifurcation in these cases corresponds to a change in the topological structure of a phase portrait as a parameter of the dynamical system is varied.
 Full characterization of the location and number of limit cycles in two-dimensional, polynomial systems is a challenging open problem in dynamical systems theory (Hilbert, 2019; Smale, 2000).
 Early theoretical results guaranteeing the existence or absence of a limit cycle in two-dimensional autonomous systems are the theorems of Poincaré-Bendixson and Bendixson-Dulac, respectively.
-Further guarantees are possible when the system can be reduced to the form of a so-called Liénard equation.
+Further guarantees are possible, for example, when the system can be reduced to the form of a so-called Liénard equation.
 None of these results, however, is constructive or easy to use in practice.
-Center manifold theory (Carr, 2012), for its part, analyzes dynamics on representative subspaces where systems reduce to normal forms having known bifurcation behaviors, though this method requires knowledge of the underlying governing equation.
+Center manifold theory (Carr, 2012), for its part, analyzes dynamics on representative subspaces where systems reduce to normal forms having known bifurcation behaviors, though this method requires knowledge of the underlying governing equations.
 The same limitation applies to Conley-Morse theory, which represents dynamics as a graph whose nodes correspond to long-term behaviors (Conley, 1978).
-Importantly, none of these approaches is adapted to real data, where equations are unknown and system measurements can be sparse, incomplete and noisy.
-Data-driven methods for detecting limit cycles Data-driven methods fall into non-topological and topological camps.
-In the former camp, several approaches use compressed sensing of trajectory data to reduce noise, followed by identification of system parameters and associated bifurcation diagrams (Wang et al., 2011; 2016), though this method requires inferring governing equations for each new sample.
+Importantly, none of these approaches is easily adapted to real data, where equations are in general unknown and system measurements can be sparse, incomplete and noisy.
+Data-driven methods for detecting limit cycles.
+Data-driven methods fall into non-topological and topological classes.
+In the former class, several approaches use compressed sensing of trajectory data to reduce noise, followed by identification of system parameters and associated bifurcation diagrams (Wang et al., 2011; 2016), though this approach requires inferring governing equations for each new sample.
 Other work sought to reduce vector field measurements to a Conley-Morse graph (Chen et al., 2007b; a) for which a database over parameter regimes can be constructed via combinatorial dynamics exploration and graph equivalence comparisons (Arai et al., 2009).
@@
 On the topological front, there is a growing body of work using persistent homology to detect periodicity in time series data (Ravishanker & Chen, 2021).
-For example, Perea & Harer (2015) extracted persistence barcodes from delay embeddings of time series to measure periodicity, although this process is very computationally expensive and there is no standard approach to classifying barcodes.
+For example, Perea & Harer (2015) extracted persistence barcodes from delay embeddings of time series data to measure periodicity, although this process is computationally expensive and there is no standard approach to classifying barcodes.
 Persistent homology approaches often assume that the underlying bifurcation parameter is known (Tymochko et al., 2020), which greatly simplifies the problem at the cost of realism.
@@
 More recent work from Ricci et al.
-(2022) uses a data-driven approach called phase2vec to learn manifolds of dynamical systems, leading to dynamical embeddings which encode qualitative properties of the underlying systems, though it is not clear if this method can detect or distinguish dynamical regimes and their associated bifurcations per se.
+(2022) uses a data-driven approach called phase2vec to learn manifolds of dynamical systems, leading to dynamical embeddings which encode qualitative properties of the underlying systems, though it is not tailored for the detection of distinct dynamical regimes and their associated bifurcations per se.
 Diffeomorphic augmentations for generalization.
-Diffeomorphic transformations which smoothly preserve topology in the context of dynamical systems have been used principally in computer vision.
+Diffeomorphic transformations which preserve topology in the context of dynamical systems have been used principally in computer vision.
 In order to train a classifier of a limited image set of biological entities such as diatoms, Vallez et al.
-(2022) used diffeomorphic augmentations to expand its train data, however, since the labels of these data were not necessarily preserved by these transformations, a manual post-filtering was required to maintain realism.
+(2022) used diffeomorphic augmentations to expand its train data, however, since the labels of these data were not necessarily preserved by these transformations, a manual post-filtering was required.
 In a second set of works concerning dynamical systems specifically, Khadivar et al.
@@
 (2021) used an autoencoder approach to reduce a complex dynamical system to a topologically conjugate (i.e., dynamically equivalent) form of discrete-time mapping.
-Here, we use diffeomorphisms to curate a set of consistently labeled dynamical augmentations, see Appendix 7.1.1.
+Here, we use diffeomorphisms to curate a set of consistently labeled dynamical augmentations.
 3 Methodology
-Topological equivalence in dynamical systems Let $\dot{x}=f(x;\theta)$ for $x\in\mathbb{R}^{2}$ be a parameterized dynamical system with parameter vector $\theta\in\Theta$.
+Topological equivalence in dynamical systems.
+Let $\dot{x}=f(x;\theta)$ for $x\in\mathbb{R}^{2}$ be a parameterized dynamical system with parameter vector $\theta\in\Theta$.
 A solution to this equation, $F(x_{0},t)$, with an initial condition, $x(t=0)=x_{0}$, is called a trajectory through $x_{0}$.
@@
 However, these properties may change as $\theta$ changes; i.e.
-the system undergoes a bifurcation.
+when the system undergoes a bifurcation.
 For example, in a system that undergoes a supercritical Hopf bifurcation, there exists a parameter $\theta$ such that below a certain threshold of $\theta$ the system exhibits a stable fixed point, and above it, the fixed point loses its stability and a stable limit cycle emerges.
-Dynamical equivalence and topological augmentation
+See Appendix 7.1 for definitions.
+Dynamical equivalence and topological augmentations.
 In what follows, we will consider autonomous systems with either a single stable fixed point (i.e., equilibrium) or a stable limit cycle with an unstable fixed point in its interior.
-We refer to systems, $f$, in one of these two categories as belonging to the same dynamical class, denoted $y(f)$, and we denote the fixed point and limit cycle classes considered here the “pre-Hopf” and “post-Hopf” classes, respectively.
+We term these two regimes as dynamical classes, denoted $y(f)$, and we denote the fixed point and limit cycle classes considered here the “pre-Hopf” and “post-Hopf” classes, respectively.
 As discussed above, these classes are invariant to topology-preserving deformations of the state space.
 Here, we will use diffeomorphisms, $h\in H$ which induce an equivalence relation on dynamical systems in a standard and well-studied way: we have $y(f)=y(f\circ h)$ for all $h\in H$ so that a given $H$ defines an equivalence relation which partitions the space of dynamical systems into its dynamical classes.
-We refer to two systems, $f,g$ such that $g=f\circ h$ as being dynamically equivalent  [^2] We choose the neutral term “dynamical equivalence” to distinguish our notion from the similar notions of topological conjugacy and topological equivalence (Kuznetsov, 1998).
-See 7.1.1 for details..
+We refer to two parameterized systems, $f,g$ such that $g=f\circ h$ as being dynamically equivalent  [^2] We choose the neutral term “dynamical equivalence” to distinguish our notion from the similar notions of topological conjugacy and topological equivalence (Kuznetsov, 1998).
+See Appendix 7.1.2 for details..
 In this light, classification becomes a problem of learning representations of dynamical systems which are invariant to all $h\in H$ but selective for the underlying dynamical class.
@@
 Our training data is constructed by first acquiring planar vector fields, $\{f(x,\theta)\in\mathbb{R}^{2}\}$ evaluated at all points, $x$, on a fixed, $64\times 64$ lattice.
-Parameters, $\theta$, are sampled uniformly on $a\in[-0.5,0.5],\omega\in[-1,1]$, and each system is assigned a label according to whether it is in the pre- or post-Hopf dynamical class.
-Then, the dataset is augmented with expressive diffeomorphisms generated using monotonic rational-quadratic splines.
+Parameters, $\theta$, are sampled uniformly on $a\in[-0.5,0.5],\omega\in[-1,1]$, and each system is assigned a label according to whether it is in the pre- or post-Hopf regime.
+Then, the dataset is augmented with diffeomorphisms generated using monotonic rational-quadratic splines.
 In particular, we use the spline construction from Durkan et al.
@@
 Because of our augmentation scheme, this classifier must contend with data which, though radically different in appearance, belong to the same dynamical class, encouraging the classifier to learn the true dynamical invariances underlying the Hopf bifurcation.
-Training and architecture details are described in the Appendix.
+Training and architecture details are described in Appendix 7.2.2.
 4 Results
@@
 We evaluate our model and its baseline competitors on a diverse, synthetic dataset of noisy (zero-mean gaussian), real-world systems drawn from across the sciences.
-Test datasets include the simple harmonic oscillator (“SO”), its warped variant (“Augmented SO”), a system undergoing a supercritical Hopf bifurcation as described in (“Supercritical Hopf”), examples based on Liénard equations (“Liénard Polynomial” and “Liénard Sigmoid”), and systems used to model oscillatory phenomena in nature and engineering applications, namely the van der Pol system, used for modeling electrical circuits in early radios (“Van der Pol”), the Belousov-Zhabotinsky model, a model of chemical oscillations (“BZ reaction”), and the Selḱov oscillator, used in modeling glycolitic cycles (“Selḱov”), see Appendix 7.2 and Strogatz (1994).
-Our baselines comprise several competing feature types: (1) “Critical Points”, a heuristic based on critical points identification as derived from Helman & Hesselink (1989; 1991) algorithm, (2) “Lyapunov” which distinguishes pre- and post-Hopf states according to the value of the Lyapunov exponent of a time series integrated from the vector field, (3) “Phase2vec”, vector field embeddings from Ricci et al.
-(2022), a state-of-the-art deep-learning method for dynamical systems embedding, and (4) “Autoencoder”, latent features of of a convolutional vector field autoencoder (the same as that used in phase2vec training), see Appendix 7.3.
+Test datasets include the simple harmonic oscillator (“SO”), its warped variant (“Augmented SO”), a system undergoing a supercritical Hopf bifurcation as described in (Strogatz, 1994)(“Supercritical Hopf”), examples based on Liénard equations (“Liénard Polynomial” and “Liénard Sigmoid”), and systems used to model oscillatory phenomena in nature and engineering applications, namely the van der Pol system, used for modeling electrical circuits in early radios (“Van der Pol”), the Belousov-Zhabotinsky model, a model of chemical oscillations (“BZ reaction”), and the Selḱov oscillator, used in modeling glycolitic cycles (“Selḱov”), see Appendix 7.3 and Strogatz (1994).
+Our baselines comprise: (1) “Critical Points”, a heuristic based on critical points identification as derived from Helman & Hesselink (1989; 1991) algorithm, (2) “Lyapunov” which distinguishes pre- and post-Hopf classes according to the value of the Lyapunov exponent of a time series integrated from the vector field, (3) “Phase2vec”, vector field embeddings from Ricci et al.
+(2022), a state-of-the-art deep-learning method for dynamical systems embedding, and (4) “Autoencoder”, latent features of a convolutional vector field autoencoder, see Appendix 7.4.
 As with our method, for all baselines except Critical Points, classifications are produced by a linear classifier trained on features from augmented simple harmonic oscillator data (Augmented SO) in order to discriminate between point and cyclic attractors.
 The accuracy average and standard deviation for our model and the trained baselines are computed over 50 training runs.
-We find that our framework generalizes across a wide range of dynamical systems and classifies distinct dynamical regimes in noisy (zero-mean gaussian) conditions better on average than baseline methods (see Table A4, Table 1 and examples at Fig.
+We find that our framework generalizes across a wide range of dynamical systems and classifies distinct dynamical regimes in noisy (zero-mean gaussian) conditions better on average than baseline methods (see Table 1, Table A4, and examples in Fig.
 A8).
@@
 Both with and without noise, Phase2vec extends mainly to SO and Supercritical Hopf.
-Autoencoder obtains competitive results, however, it exhibits little generalization to Liénard Polynomial system, and essentially no generalization to Selḱov system.
+Autoencoder obtains competitive results but exhibits little generalization to Liénard Polynomial system, and essentially no generalization to Selḱov system.
 We also test resilience to noise in the Augmented SO data (see Table 2).
@@
 Autoencoder 0.86$\pm$0.01 0.87$\pm$0.01 0.84$\pm$0.02 0.8$\pm$0.04 0.76$\pm$0.05 0.72$\pm$0.05 0.69$\pm$0.06 0.68$\pm$0.06 0.65$\pm$0.06 0.64$\pm$0.05 0.62$\pm$0.05
-In the Appendix 7.3, we report the results of another baseline, “Parameters” which trains a linear classifier based on a flattened vector of coefficients of a degree three polynomial fit to each vector field with least squares, a common method for bifurcation analysis, although this resulted in a poor average accuracy of 53%.
+In Appendix 7.4, we report the results of another baseline, “Parameters” which trains a linear classifier based on a flattened vector of coefficients of a degree three polynomial fit to each vector field with least squares, a common method for bifurcation analysis, although this resulted in a poor average accuracy of 53%.
 Furthermore, we find that all modules in our pipeline (topological augmentation, angular representation of data, self-attention layers) contribute, on average, to accuracy in complex systems, see Appendix 7.5.
@@
 These overall suggest that our method can be geared towards recovering the bifurcation boundaries of unseen systems.
-Figure 2: Inference of bifurcation boundary.
+Figure 2: Inference of bifurcation boundaries.
 Each system is plotted at coordinates of its simulated parameter values and colored according to its average cycle prediction across 50 training re-runs.
@@
 In order to demonstrate the viability of our method for higher-dimensional systems under realistic noise conditions, we apply our topological approach to a computational model in a well-studied biological setting, the so-called repressilator model (Elowitz & Leibler, 2000).
-The repressilator system is a synthetic, genetic regulatory network that was introduced by Elowitz and Liebler into bacteria to induce oscillatory regulation of gene expression Fig.
-3.
+The repressilator system is a synthetic, genetic regulatory network that was introduced by Elowitz and Liebler into bacteria to induce oscillatory regulation of gene expression (Fig.
+3).
 The system is modeled by the mRNA and protein counts of three genes (TetR, LacI and $\lambda$ phage cI) that inhibit each other in a cyclic manner (see Appendix 7.8 for governing equations).
 Theoretically and experimentally, the existence and characteristics of oscillations have been shown to depend on numerous factors like the rates of transcription, translation, and degradation, as well as intrinsic and extrinsic noise (Verdugo, 2018; Potvin-Trottier et al., 2016).
-Here we test our method in classifying a regime of the repressilator which has been shown to be undergoing a Hopf bifurcation, though, unlike the previous complex simulated systems, it spans six dimensions and is constructed from noisy cellular time-series.
-To generate 2D vector field data, we consider trajectories of cellular gene expression measured by protein levels of TetR and LacI (as in Fig.
+Here we test our method in classifying a regime of the repressilator which has been shown to undergo a Hopf bifurcation, though, unlike the previous complex simulated systems, it spans six dimensions and is constructed from noisy cellular time-series.
+To generate 2D vector field data, we consider trajectories of cellular gene expression measured by protein levels of TetR and LacI (Fig.
 3, middle).
 Different instances of simulations share the initial condition but vary by the rate of transcription, $\alpha$, and by the ratio of protein and mRNA degradation rates, $\beta$.
-We sample 100 cell states along the temporal trajectory and add Gaussian noise of $\sigma=0.5$ to the gene expression states (to protein and mRNA levels) to model intrinsic noise.
-For each set of measurements, we interpolate an input 64x64 vector field, based on their velocities across TetR-LacI protein phase space.
-For further details of the data generation process see Appendix 7.7.
+We sample 100 cell states along the temporal trajectory and add Gaussian noise of $\sigma=0.5$ to the gene expression states (to protein and mRNA levels).
+For each set of measurements, we interpolate an input 64x64 vector field, based on their velocities across TetR-LacI protein phase space, see examples at Fig.
+A14 and Appendix 7.8 for further details.
 We predict for each vector field its dynamical class, achieving 87% classification accuracy and proximate bifurcation boundary to the analytically-derived boundary (Fig.
@@
 A13 we show prototypes of trajectories across the $\alpha-\beta$ parameter space, where we see that where we fail to classify a point attractor, trajectories of cells loop closely in the TetR-LacI protein plane and generate data samples that highly resemble trajectories with a periodic attractor rather than with a point attractor.
-Figure 3: Oscillations in repressilator system.
+Figure 3: Oscillations in a repressilator system.
 Example of oscillations in the repressilator system viewed as a time-series (left) or projected onto TetR-LacI protein plain (middle).
@@
 We show that our model can detect the qualitative change in the flow of cells in gene expression space as they mature and differentiate out of their precursor cell state, moving from a dynamical regime characterized by a periodic behavior to that characterized by fixed points.
-We focus on the process of pancreas endocrinogenesis where Ductal and Ngn3+ precursor cells rapidly proliferate to give rise to functional cell types Alpha, Beta, Delta and Epsilon (Bastidas-Ponce et al., 2019).
-We examine a single-cell RNA-sequencing dataset where the expression of over 2,000 genes are measured per cell, for a total of 3,600 cells.
+We focus on the process of pancreas endocrinogenesis where Ductal and Ngn3+ precursor cells rapidly proliferate to give rise to functional cell types Alpha, Beta, Delta and Epsilon (Bastidas-Ponce et al., 2019); In that study, a single-cell RNA-sequencing dataset was collected where the expression of over 2,000 genes are measured per cell, for a total of 3,600 cells (Fig.
+4, left).
 While the RNA expression profile, providing a gene expression vector per cell, provides for each cell its location within gene expression space, RNA velocity (Bergen et al., 2021; Gorin et al., 2022) can be used to compute the effective velocity of each cell (based on differences in spliced and unspliced RNA counts).
-We use scVelo (Bergen et al., 2020) to infer cellular velocities in the high-dimensional gene expression space and to project them to a 2D UMAP representation (Fig.
-4, left).
+We use scVelo (Bergen et al., 2020) to infer cellular velocities in the high-dimensional gene expression space and to project them to a 2D UMAP representation.
 Simultaneously, we also compute for each cell a ‘cell-cycle score’, which captures the extent of expression of cell-cycle genes (in phase S and in grouped G2 and M phases, Fig.
@@
 To generate multiple frames from the trajectory, we randomly split cells of each cell type into groups of at least 50 cells resulting in 69 (sparse, noisy) vector fields with corresponding cell-cycle score (Fig.
-4, middle, see details in Appendix 7.8).
+4, middle, Appendix 7.9).
 We then used our model (still trained solely on simulated augmented simple oscillator data) to characterize the dynamic regimes of the cellular behavior.
-Our model can successfully identify proliferative, cyclic behavior within the single-cell data, as reflected by 94% accuracy in predicting a cyclic behavior matching the high cell cycle score (Fig.
-4, right), exceeding the prediction accuracy reached by alternative baselines (Critical Points: 54%, Lyapunov: 49%, Phase2vec: 30%, Autoencoder: 30%).
-As opposed to the cell-cycle score, our model does not use any prior knowledge regarding cell-cycle related genes, and is only based on the structural features of the cellular vector fields, and therefore can also be employed when such prior knowledge is not available.
+Our model can successfully identify proliferative, cyclic behavior within the single-cell data, as reflected by 94% accuracy in predicting a cyclic behavior matching high cell cycle scores (Fig.
+4, right), substantially exceeding the prediction accuracy reached by alternative baselines (Critical Points: 54%, Lyapunov: 49%, Phase2vec: 30%, Autoencoder: 30%).
+As opposed to the cell-cycle score, our model does not use any prior knowledge regarding cell-cycle related genes, and is only based on the structural features of the cellular vector fields, and therefore can also be employed when such prior knowledge is unavailable.
 Figure 4: Proliferation-to-differentiation transition in pancreas development.
@@
 Colors distinguish between different cell types.
-(middle) Cell-cycle score based on the fraction of cells with high S-phase gene expression (with S-score $>$ 0.4; see Appendix 7.8).
+(middle) Cell-cycle score based on the fraction of cells with high S-phase gene expression (with S-score $>$ 0.4).
 (right) Predicted cell-cycle score computed as the average classification prediction across 50 training re-runs.
+See details in Appendix 7.9.
 5 Limitations
@@
 Scaling up to much higher dimensions in these contexts presents technical challenges like the exponential growth of input size with dimension, as well as theoretical challenges such as the lack of structural stability in higher dimensions and the introduction of chaotic regimes.
-Since inherently low-dimensional dynamics often underlie high-dimensional measurements, especially in biology (Xiong & Garfinkel, 2023; Khona & Fiete, 2022), high dimensions can be mitigated by hardwired, or optimized dimensionality reduction methods as we deploy for the six-dimensional repressilator and for analyzing over 2,000 gene expression dynamics of the pancreas single-cell dataset.
+Since inherently low-dimensional dynamics often underlie high-dimensional measurements, especially in biology (Xiong & Garfinkel, 2023; Khona & Fiete, 2022), high dimensions can be mitigated by hardwired, or optimized dimensionality reduction methods as we deploy for the six-dimensional repressilator and for analyzing over 2,000 gene expression (dimensional) dynamics of the pancreas single-cell dataset.
 Additionally, we used a supervised, multi-class setup so that the classifier can predict that both a limit cycle and a fixed point co-exist.
@@
 To accommodate more versatile invariances and bifurcations, future work could extend our framework to the unsupervised case.
-For example, given that our data augmentations produce samples which are in the same dynamical equivalence class, a contrastive loss could be used to shape the feature space so that bifurcation classes are represented implicitly during training.
+For example, given that our data augmentations produce samples which are in the same dynamical equivalence class, a contrastive loss could be used to shape the feature space so that dynamical classes are represented implicitly during training.
 Or, while challenging, following Skufca & Bollt (2008), a purely unsupervised approach could be used to explicitly construct warpings between systems in order to characterize equivalence classes.